OpenStudy (anonymous):

Solve for x 5(x + 1) + 3(3 – x) = 0

OpenStudy (anonymous):

5x +5 +9 - 3x = 0 2x +14 = 0 2x = -14 x = -7

OpenStudy (aroub):

5x+5+9-3x=0 2x+14=0 2x=-14 x=-7

OpenStudy (anonymous):

thnx 15starjordan!

OpenStudy (anonymous):

and aroub

OpenStudy (anonymous):

-7